# Accounting for Focus Ambiguity in Visual Questions

### v1

@@ 10.1 Recognizing Questions with Focus Ambiguity
Figure 18: Qualitative results in end-to-end localization and multi-step localization illustrate various challenges in the localization task.
Qualitative Results.
@@ 10.2 Locating All Plausible Regions of Focus
Both the end-to-end approach and the multi-step approach perform poorly in question focus localization. From the quantitative results in Section 4.2, we discover that the models struggle to localize regions in ambiguous questions (i.e., multiple groundings), regions in PACO-LVIS, especially those that are parts instead of objects in the dataset, and regions that don’t align with answer grounding. To highlight these challenges, we provide qualitative results in Figure 18. Example 1 illustrates a case where both the end-to-end approach and the multi-step approach locate only the largest plausible focus region and miss the other regions. Example 2 illustrates a case where the end-to-end approach grounds both the question region and answer regions, while multi-step approach generates only the answers in step 1 and thus only grounds answers. Example 3 illustrates a case where the end-to-end approach localizes the question region relatively well, and step 1 in the multi-step approach generates a correct description but completely misses the described area in step 2. From these examples, we discover that both approaches demonstrate confusion between question groundings and answer groundings. Also, in multi-step approach, challenges can occur in both steps.
Figure 19: Analysis on the results of step 1 in multi-step approach for locating question focus regions. (a) Distribution of the predicted number of descriptions in the first step by the number of question focus regions in ground truth. (b) Example of multiple descriptions generated in step 1 and their grounding results compared to results of the end-to-end approach.
Analysis on Generated Descriptions in Multi-Step Approach.
For the question focus grounding task, we introduce the multi-step approach to compensate for the poor performance of GLaMM in generating multiple regions in one answer. However, we did not observe an obvious improvement in the overall performance of the multi-step approach. We further break down the results of the first step, describe, and the second step, localize, to highlight the bottleneck of the task. In Figure 19(a), we present the distribution of the predicted number of descriptions in the first step by the number of question focus regions in ground truth. We discovered that the overall segmentation performance is poor, and also the number of descriptions generated in the first step doesn’t have a strong correlation to the number of ground truth regions. In Figure 19(b), we observe that the number of described question focus regions increased in the multi-step approach compared to the end-to-end approach. However, we can still see that the model suffers from not generating a clear description for every region. We suspect that the poor performance in the first step, describe, might impact the performance of the multi-step approach and thus reduce the improvement in its overall performance.

### v2

@@ 9 Model Analysis
9.1 Model Details
We used for recognition task ChatGPT-4o, InternVL2-Llama3-76B, Qwen2.5-VL-3B-Instruct, Qwen2.5-VL-7B-Instruct, Qwen2.5-VL-72B-Instruct,and Molmo-7B-D-0924. The models we used for localization task are Molmo-7B-D-0924, GLaMM-FullScope, ChatGPT-4o, and SAM-ViT-h-4B.
9.2 Prompting Methods
Ablation Pilot Study for Prompts.
To settle on the basic prompt (ZS), we first tested four variants of focus ambiguity definition on the GPT-4o model in zero-shot settings on all visual questions in VizWiz-VQA and VQAv2 for the recognition task as a pilot study. For these prompts, we substituted the text by modifying the definition of focus ambiguity in “You are a helpful assistant. You will be given an image and a question about the image. Your task is to predict whether the question is ambiguous or unambiguous based on the given image and the definition of focus ambiguity. Focus ambiguity in a visual question occurs when…(definition of focus ambiguity). Please only output “ambiguous” or “unambiguous”.”. The four different definitions are: (1) “if under-specified language can be grounded to multiple regions in the image” (F1: 78, Acc: 76), (2) “when the question does not clearly indicate which part of the image it specifies” (F1: 76, Acc:74), (3) “if the question refers to multiple regions in the image.” (F1: 74, Acc:73) (4) and “if there are multiple regions in the image each satisfying the question’s constraints” (F1: 77, Acc: 75). The phrasing “if under-specified language can be grounded to multiple regions in the image” proved to be the most effective and thus we selected this definition.
We then tested whether introducing the term “focus ambiguity” to the model in the prompting as opposed to simply providing the conditions “An visual question is ambiguous if …. It is unambiguous if…”. The results show a boost to an F1 score of 80 for the AnswerTherapy source. Therefore, we used this last format. The final prompt for zero-shot setting is shown in Figure 18.
{promptbox}
“You are a helpful assistant. You will be given an image and a question about the image. Your task is to predict whether the question is ambiguous or unambiguous based on the given image.
A question is ambiguous if under-specified language in questions can be grounded to multiple regions in the images, leading to uncertainty about the intended focus. Specifically, ambiguity arises when multiple regions in the image could each satisfy all the constraints of the question, but the question fails to clarify which one it refers to (e.g., by not using the plural form or failing to specify a region).
A question is unambiguous if it clearly identifies or specifies the part of the image it refers to, leaving no uncertainty about the intended focus. If the image contains only one object or region that satisfy all the constraints of the question, and there is no possibility of multiple interpretations due to its simplicity, the question is considered unambiguous, even if the question lacks specificity or the phrasing is vague (e.g., ‘What is the color of the object?’ when there is only one object shown in the image); If the question explicitly refers to multiple regions using the plural form or clear descriptors, the question is also considered unambiguous.
Please only output “ambiguous” or “unambiguous”.
Figure 18: The prompt for Zero-Shot setting.
{promptbox}
“(omitted, same as zero-shot setting…)
Please think step by step first, and then predict whether it is ambiguous or not.
Please only output “ambiguous” or “unambiguous”.
Figure 19: The prompt for ZS-COT setting.
{promptbox}
“(omitted, same as zero-shot setting…)
Please follow these steps to think step by step first, and then predict whether it is ambiguous or not:
1. Understand the Image: Carefully analyze and fully comprehend the given image.
2. Understand the question: Carefully analyze and fully comprehend the question.
3. Find the regions: Find all regions in the image that could each satisfy all the constraints and contain all the necessary information to answer the question.
4. Predict: Predict whether the question is ambiguous or not. If the there are multiple regions found and the question does not use plural form, it is an ambiguous question. Otherwise, it is an unambiguous question.
Please only output “ambiguous” or “unambiguous”.
Figure 20: The prompt for ZS-ECOT setting.
{promptbox}
“(omitted, same as zero-shot setting…)
Here are two examples with the images described for you:
Example 1:
Question: What color is the kite?
Image description: The image shows a group of people in a park with picnic tables. One table has some food on it and there are two women sitting at the table. Other people are standing. One young boy is holding a blue kite in his hand, another child is picking up a blue kite from the ground, and one man is running and flying a yellow kite.
Prediction: ambiguous
Example 2:
Question: Where is the man holding the apple?
Image description: The image shows a man wearing a blue shirt sitting at a dining table and a little girl sitting on the floor beside him. The setting is likely a dining room. There is a window behind the dining table with brown curtains. The man is holding an apple in his hand. The little girl is holding a toy in her hand. A brown dog is laying on the floor looking at the girl.
Prediction: unambiguous
Below is the task for you to make prediction based on the given image. Please only output “ambiguous” or “unambiguous”.
Figure 21: The prompt for FS setting.
{promptbox}
“(omitted, same as zero-shot setting…)
Please follow these steps to think step by step first, and then predict whether it is ambiguous or not:
1. Understand the Image: Carefully analyze and fully comprehend the given image.
2. Understand the question: Carefully analyze and fully comprehend the question.
3. Find the regions: Find all regions in the image that could each satisfy all the constraints and contain all the necessary information to answer the question.
4. Predict: Predict whether the question is ambiguous or not. If there are multiple regions found and the question does not use plural form, it is an ambiguous question. Otherwise, it is an unambiguous question. Please only output “ambiguous” or “unambiguous”.
Here are two examples with the images described for you:
Example 1 [INSERT SAME EXAMPLE 1 AS FS SETTING]
Step 1. Understand the Image: Understand the given image as described.
Step 2. Understand the question: The question is about the color of the object, the kite, in the image.
Step 3. Find the regions: The three regions of the three kites are the regions that can each satisfy all the constraints and contain all the necessary information to answer the question.
Step 4. Predict: There are three regions found — multiple regions make this question ambiguous. Therefore, the prediction is “ambiguous”.
Example 2: [INSERT SAME EXAMPLE 2 AS FS SETTING]
Step 1. Understand the Image: Understand the given image as described.
Step 2. Understand the question: The question is about where the man holding the apple is.
Step 3. Find the regions: In the image we can only find one man holding the apple, which satisfies all the constraints and contains all the necessary information to answer the question.
Step 4. Predict: There is only one region found. Therefore, the question is “unambiguous”.
Below is the task for you to make prediction based on the given image. Please follow the steps to predict whether the question is ambiguous or not. Please only output “ambiguous” or “unambiguous”.
Figure 22: The prompt for FS-ECoT setting.
Details of Prompts for Each Experiment.
In model benchmarking, all models are tested with five prompting methods, zero-shot (ZS), zero-shot chain of thought (ZS-CoT), zero-shot enhanced chain of thought (ZS-ECoT), few-shot (FS), and few-shot enhanced chain of thought (FS-CoT). These are illustrated in Figure 18192021, and 22.
For focus ambiguity recognition, we prompted GPT-4o, InternVL2, Qwen2, and Molmo using these five prompts, respectively, to acquire classification results.
For end-to-end region localization, we prompted GLaMM by adjusting these five prompts to “segment each of the regions in one mask.” For example, for step 4 of the example 1 of FS-ECoT, we adjusted it to “Predict segmentation: Predict three segmentation masks separately - segmentation of the blue kite, segmentation of the red kite, and segmentation of the yellow kite.”
For ChatGPT-4o+GLaMM localization, we prompted GPT-4o using these five methods to acquire descriptions of the regions. We add additional formatting constraints let GPT-4o only “Use a word or a phrase to describe each region. Please only output the descriptions of these regions and use commas to divide descriptions of regions if there are multiple regions.” We then prompt GLaMM using the each generated description embedded in “Can you segment {description}?” to acquire segmentation masks.
For Molmo+SAM localization, we used the same five prompts with the only difference being replacing the “segment” as “point” for Molmo generate points. The generated points’ coordinates are fed into SAM to generate the segmentations.
@@ 9.2 Prompting Methods
Discussion about Prompt Strategies.
As shown in the main paper in section 4.1, when comparing different prompt types, we found that across all models, ZS-CoT and ZS-ECoT consistently lead to better performance. This is because, CoT is better than without CoT (not surprising); and surprisingly, ZS is better than FS across all models, except for Molmo (7B), in our setting.
We attribute this unexpected outcome to three interconnected factors: The few shot setting has (1) too limited (only two shots) and (2) challenging examples, (3) without additional explanations along with the examples. We intentionally selected two challenging examples for FS, including one unambiguous question with answer ambiguity, as shown in Figure 21. The original intention was to encourage models to differentiate clearly between question and answer ambiguity, similar to how contrastive learning and humans benefit from hard samples. However, with only two shots, it turned out to only confuse models and led to overfitting towards ambiguous cases, possibly because the unambiguous case is too hard for model to understand without first learning from simpler, more basic cases, without additional explanation. We attribute this to the model exhibiting overfitting tendencies, with only the ambiguous case understood, evidenced by a significantly increased positive rate in FS for all models.
In contrast, FS-ECoT includes step-by-step explanations, significantly boosting performance with the same two hard examples. Given that most models are sensitive to prompt types, as shown in Table 2 in the main paper, we encourage future research to further explore prompt strategies, starting with basic and simple examples, and explore in-context fine-tuning/expanding context windows/integrating multi-image inputs to enable a richer set of few-shot examples, and provide explicit reasoning steps (e.g., ECoT) to help models to understand examples.
@@ 9.3 Recognizing Questions with Focus Ambiguity
Different Sizes of Qwen2.5-VL.
Model size ZS ZS-CoT ZS-ECoT FS FS-ECoT
72B 58.7 60.5 64.7 57.7 64.7
7B 57.2 63.8 65.5 53.6 59.0
3B 44.6 44.6 45.0 44.3 44.3
Table 5: Performance of Qwen2.5-VL model series in three model sizes, 72B, 7B, and 3B, presented in accuracy.
We benchmarked Qwen2.5-VL across model sizes ranging from 72B to 3B, as shown in Table 5. The 72B model achieves the best performance in ZS, FS, and FS-ECoT settings, while the 7B model performs best in ZS-CoT and ZS-ECoT. Notably, the 7B model with ZS-ECoT yields the highest overall performance across all sizes and prompting strategies, so we report its results in the main paper. In contrast, the 3B model consistently underperforms, which we attribute to its limited reasoning capabilities due to its significantly smaller LLM size.
Effect of Fine-tuning on Qwen2.5-VL.
To understand whether fine-tuning on our dataset can improve recognizing question ambiguity, we fine-tuned Qwen2.5-7B improves accuracy by 1.6 percentage points to 58.8% (ZS), matching the larger Qwen2.5-72B (ZS). While we expect similar boosts for other models, overall performance remains low.
Large reasoning model.
we conducted an ablation study with ChatGPT-o3 ($\gtrsim$ 200B). The results, (Table 6), reinforce that existing models struggle to recognize question ambiguity.
Ablation Study on Input Types for ChatGPT-4o.
We conducted two ablation studies to assess the impact of different input types on ChatGPT-4o’s ability to recognize questions with focus ambiguity.
Question-only input. Given that existing models often exhibit strong language priors and may overlook visual input, a common issue in the vision-and-language domain, we evaluated ChatGPT-4o using only the question text. As shown in Table 6, performance drops significantly across all prompts, with decreases ranging from 15.2% to 25.9%, highlighting the critical role of image input.
Caption as image substitute. In our few-shot setting, example inputs consist of a question paired with a textual description of an image, while the actual task involves answering a question given an image, due to the model’s current limitation in processing multiple images. To explore this modality mismatch, we benchmarked few-shot ChatGPT-4o using its own generated captions in place of image inputs at inference time. The results are comparable (Table 6); for example, few-shot accuracy increased by 1.3%, while FS-ECoT decreased by 2.1%.
ZS ZS-CoT ZS-ECoT FS FS-ECoT
ChatGPT-o3 63.0 66.6 68.1 64.7 70.9
ChatGPT-4o (Q only) 44.2 45.9 42.5 44.8 44.9
ChatGPT-4o (ImgCap) - - - 61.3 62.8
Table 6: Recognition (accuracy) results.
Figure 24: Qualitative results in GLaMM and ChatGPT-4o+GLaMM, illustrating various challenges in the localization task.
Qualitative Results
@@
Success Case of GPT-4o.
We investigate the best performing model, GPT-4o, for the recognition task. We attribute GPT-4o’s advantage to its bigger size compared to other models as well as its extensive training with human alignment and real-world feedback, where ambiguity is likely common. To further explore GPT-4o’s success, we present qualitative results in Figure 25, showing samples where GPT-4o succeeds while other models fail to recognize the focus ambiguity. We found that GPT-4o is superior in identifying unambiguous samples compared to other models, especially when other models also overlook constraints in the question; e.g., for “What color is the largest microwave?” and when the question uses a plural form, e.g., “What items are on the plates?”
Figure 25: Qualitative results for examples where GPT-4o succeeds while other models fail.
Failure Case of GPT-4o.
We then qualitatively review GPT-4o’s prediction errors and found it often struggles when there is question ambiguity while the image prominently features a salient entity, as exemplified in Figure 26(a); the model tends to interpret the scenario as unambiguous, ignoring other plausible entities that are less prominent. The model can also mistakenly deem unambiguous questions as ambiguous by overlooking constraints in the question; e.g., for “What is the title of the book with a blue cover in the center of the image?” only one of multiple books satisfies the constraint of having a blue cover. Other cases include (1) the question uses a plural form (2) the question is unambiguous but the answer involves ambiguity, the model often provides ambiguous predictions. For example, in the question “What is on the table?”, if there is only one table but multiple types of food on it, as shown in Figure 26(b) in the left image, GPT-4o fails by predicting it as ambiguous.
Figure 26: Qualitative results for GPT-4o, where (a) presents ambiguous question but GPT-4o predicted it as unambiguous and (b) presents unambiguous question but GPT-4o predicted it as ambiguous.
Undecided Rate.
The undecided rate indicates how often a model abstains to make a focus ambiguity prediction. We expected this rate to be inversely correlated with model performance, but surprisingly, this was not the case. For example, CoT, with better performance, often results in higher undecided rates. Another example is the worst-performing model, Qwen (3B), has the lowest undecided rate.
We suspect that models with stronger reasoning skills might realize and acknowledge its limitation rather than being over-confident and hallucinating; i.e. known unknowns vs. unknown unknowns. From qualitative results, we observe that undecided responses typically include phrases like “I’m unable to fulfill this request” and ”I need more information”.
@@ 9.4 Locating All Plausible Regions of Focus
The GLaMM approach, ChatGPT+GLaMM, and Molmo+SAM all perform poorly in question focus localization. From the quantitative results in Section 4.2, we discover that the models struggle to localize regions in ambiguous questions (i.e., multiple groundings); regions in PACO-LVIS, especially those that are parts instead of objects in the dataset; and regions that don’t match answer groundings. To highlight these challenges and supplement the qualitative results for Molmo+SAM presented in the main paper, we provide additional qualitative results in Figure 24 for the other two localization methods. Example 1 illustrates a case where both the GLaMM approach and the ChatGPT4o+GLaMM approach locate only the largest plausible focus region and miss the other regions. Example 2 illustrates a case where the GLaMM approach grounds both the question region and answer regions, while ChatGPT4o+GLaMM generates only the answers in step 1 and thus only grounds answers. Example 3 illustrates a case where the GLaMM approach localizes the question region relatively well, and step 1 in the ChatGPT4o+GLaMM generates a correct description but completely misses the described area in step 2. From these examples, we discover that both approaches demonstrate confusion between question groundings and answer groundings. Also, in ChatGPT+GLaMM, challenges can occur in both steps.
Figure 27: Analysis on the results of the prediction from GPT-4o in the GPT-4o + GLaMM approach for locating question focus regions. (a) Distribution of the predicted number of descriptions in the first step by the number of question focus regions in ground truth. (b) Example of multiple descriptions generated in step 1 and their grounding results compared to results of the end-to-end approach.
Analysis on Generated Descriptions in ChatGPT+GLaMM Approach.
For the question focus grounding task, we introduce the ChatGPT+GLaMM approach to compensate for the poor performance of GLaMM in generating multiple regions in one answer. However, we did not observe an obvious improvement in the overall performance of the ChatGPT+GLaMM approach. We further break down the results of the first step, describe, and the second step, localize, to highlight the bottleneck of the task. In Figure 27(a), we present the distribution of the predicted number of descriptions in the first step by the number of question focus regions in ground truth. We discovered that the overall segmentation performance is poor, and also the number of descriptions generated in the first step doesn’t have a strong correlation to the number of ground truth regions. In Figure 27(b), we observe that the number of described question focus regions increased in the ChatGPT+GLaMM approach compared to the end-to-end GLaMM approach. However, we can still see that the model suffers from not generating a clear description for every region. We suspect that the poor performance in the first step, describe, might impact the performance of the ChatGPT+GLaMM approach and thus reduce the improvement in its overall performance.